OpenStudy (anonymous):

Write a system of two equations using x and y to represent the following problem. A chemist is mixing two solutions together. The chemist wants 100 gallons when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use?

OpenStudy (diyadiya):

Let x be the amount of 24% soln and y be the amount of 50% soln The chemist wants 100 gallons when complete ,So \[x+y = 100 \tag{1}\]He is mixing a 24% solution with a 50% solution to get a 36% solution (36 % of final soln) 24/100x + 50/100y = 36/100*100 (its 36% of 100 gallons) \[0.24x+0.50y=36 \tag{2}\] Now you can solve these two equations